A random sample of 815 adults surveyed resulted in 20% preferred honey to jam/jelly on their peanut butter sandwich. Find the 99% confidence interval for the proportion of adults that prefer honey.

What is the UL, Upper limit of the confidence interval?

0.233 is not the answer... :(
Answer #1

Solution :

n = 815

= 0.200

1 - = 1 - 0.250 = 0.750

At 99% confidence level the z is ,

  = 1 - 99% = 1 - 0.99 = 0.01

/ 2 = 0.01 / 2 = 0.005

Z/2 = Z0.005 = 2.576

Margin of error = E = Z / 2 * (( * (1 - )) / n)

= 2.576 * (((0.200 * 0.800) / 815)

= 0.036

A 90 % confidence interval for population proportion p is ,

+ E

0.200 + 0.036

= 0.236

Upper limit = 0.236